OpenStudy (erinkb99):

Find the range of the function F(x)= integral from [-4, x] sqrt(16-t^2) dt

jimthompson5910 (jim_thompson5910):

hint: y = sqrt(16-x^2) is a semi-circle with radius 4

OpenStudy (erinkb99):

so [-4,4]?

jimthompson5910 (jim_thompson5910):

the integral represents the area under the curve, which in this case is the area of a part of the semi-circle or all of it

jimthompson5910 (jim_thompson5910):

what is the smallest area possible?

OpenStudy (erinkb99):

0

jimthompson5910 (jim_thompson5910):

yes, so that is the lower bound of the range

jimthompson5910 (jim_thompson5910):

what is the largest area possible? ie, what is the area of the semi-circle?

OpenStudy (erinkb99):

[0, 4] because its asking for range and that encompasses the highest point

OpenStudy (erinkb99):

right?

OpenStudy (erinkb99):

4pi

jimthompson5910 (jim_thompson5910):

area of semi-circle formula \[\Large A = \frac{\pi*r^2}{2}\]

jimthompson5910 (jim_thompson5910):

in this case, r = 4

OpenStudy (erinkb99):

so the largest area is 8pi

jimthompson5910 (jim_thompson5910):

yes, which is the upper bound of the range

jimthompson5910 (jim_thompson5910):

therefore, the range is from 0 to 8pi

OpenStudy (erinkb99):

ahhh ok

jimthompson5910 (jim_thompson5910):

in interval notation, you'd write \(\Large [0, 8\pi]\)

OpenStudy (erinkb99):

I understand. I'm posting a few more questions. Mind helping out?

jimthompson5910 (jim_thompson5910):

Sure, go ahead